Brief Exercise 13-10

Suppose these selected condensed data are taken from recent balance sheets of Bob Evans Farms (in thousands).

2022

2021

Cash

$13,100

$10,800

Accounts receivable

23,800

20,200

Inventory

30,200

30,800

Other current assets

11,500

11,909

Total current assets

$78,600

$73,709

Total current liabilities

$228,500

$299,600


Compute the current ratio for each year. (Round answers to 2 decimal places, e.g. 2.12.)

2022

2021

Current ratio:

enter the current ratio for year 2017 rounded to 2 demical places

:1 enter the current ratio for year 2016 rounded to 2 demical places

:1
Answer #1

Current ratio = current assets/current liabilities

2022 ratio = 78,600/228,500 = 0.34

2021 ratio = 73,709/299,600 = 0.25